# Misspecification uncertainties in near-deterministic regression

### v2

@@ III Analysis of the generalization error in the near-deterministic limit
for any $c>0$. Equation (15) is our second main result, a demonstration that any candidate minimizer of the generalization error must have probability mass in every POPS ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$, whilst the minimizing distribution concentrates on regions where multiple POPS intersect. As shown above, the minimum loss solution $\pi_{\mathcal{L}}({\bm{\Theta}})$ does not satisfy this requirement and thus ${\mathcal{G}}[\pi_{\mathcal{L}}]=\mathcal{O}(1/\epsilon^{2})$. Importantly, when (15) is satisfied, the predictive distribution $\rho_{\mathcal{M}}({\mathbf{Y}},{\mathbf{X}})$ covers every training data point, i.e. model predictions are guaranteed to envelope observations as $N/P\to\infty$.
III.1 POPS-constrained loss minimizing ensemble ansatz
The covering constraint (15) shows the minimizer of ${\mathcal{G}}_{0}$ must have mass in the POPS of each training point. This suggests that a representation in terms of a discrete set of models is likely to capture its key features. Selecting some set of pointwise fits $\mathcal{T}_{E}=\{{\bm{\Theta}}_{\mathbf{X}}\}_{{\mathbf{X}}\in{\mathcal{X}}}$, ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$ for each ${\mathbf{X}}\in{\mathcal{X}}$ and some weighting ${\rm w}_{E}({\mathbf{X}})>0$, $\langle{\rm w}_{E}({\mathbf{X}})\rangle_{\mathcal{X}}=1$, equation (15) is satisfied by the Gaussian ensemble ansatz
$\pi_{E}({\bm{\Theta}})\equiv\langle{\rm w}_{E}({\mathbf{X}})\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{{\mathbf{X}}},\sigma^{2}\mathbb{I})\rangle_{\mathcal{X}},$ (16)
where $\sigma$ is fixed and assumed small. Defining mass function for the ensemble which measures the number of pointwise fits ${\bm{\Theta}}_{{\mathbf{X}}^{\prime}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$
$m_{E}({\mathbf{X}})\equiv\left\langle{\rm w}_{E}({\mathbf{X}}^{\prime}){\rm m}_{\mathcal{E}}({\bm{\Theta}}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})\right\rangle_{{\mathcal{X}}},$ (17)
the generalization error of (16) reads, in the limit $\sigma\to 0$,
${\mathcal{G}}_{0}[\pi_{E}]=-\left\langle\ln|m_{E}({\mathbf{X}})|\right\rangle_{{\mathcal{X}}}.$ (18)
The above result shows to minimize ${\mathcal{G}}_{0}[\pi_{E}]$ we should select fits ${\bm{\Theta}}_{\mathbf{X}}\in\mathcal{T}_{E}$ to have maximum occupancy across all other POPS. In the worst case $\mathcal{T}_{E}\cap{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})={\bm{\Theta}}_{{\mathbf{X}}}$, $\forall{\mathbf{X}}\in{\mathcal{X}}$, ${\mathcal{G}}_{0}[\pi_{E}]$ still has a finite upper bound as $\sigma\to 0$ of
${\mathcal{G}}_{0}[\pi_{E}]\leq-\left\langle\ln|{\rm w}_{E}({\mathbf{X}})|\right\rangle_{{\mathcal{X}}}+\mathcal{S}_{{\mathcal{X}}},$ (19)
where $\mathcal{S}_{{\mathcal{X}}}=-\left\langle\ln|\rho({\mathbf{X}})|\right\rangle_{{\mathcal{X}}}$ is the entropy of the input distribution. In appendix D we show how each ${\bm{\Theta}}_{\mathbf{X}}\in\mathcal{T}_{E}$ can be determined in a variational setting at finite $\sigma$, using (16) in the PAC-Bayes loss bound (7) for the regression loss ${\mathcal{L}}[\pi_{E}]=\langle{\rm w}_{E}({\mathbf{X}})L({\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}$, which as $P/N\to 0$, $\sigma\to 0$ reduces to POPS-constrained loss minimization
${\bm{\Theta}}^{*}_{\mathbf{X}}$ $\equiv\arg\min_{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}L({\bm{\Theta}}_{\mathbf{X}}),$ (20)
$\pi^{*}_{E}({\bm{\Theta}})$ $\equiv\langle{\rm w}_{E}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{{\mathbf{X}}})\rangle_{\mathcal{X}}.$ (21)
The goal of this POPS-constrained loss-minimization is to ensure the pointwise fits will be as close as possible; for convex loss landscapes, this is measured by the Fisher information metric ${\bm{\mathcal{I}}}^{*}_{\mathcal{L}}$. By minimizing the ensemble spread, we aim to distribute an (albeit small) mass to other POPS, thereby reducing ${\mathcal{G}}_{0}$. In addition, for specified regression problems $\pi^{*}_{E}({\bm{\Theta}})$ then reduces to the true minimizer $\pi^{*}_{\mathcal{L}}({\bm{\Theta}})=\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$.
When the loss landscape is not convex additional constraints will have to be imposed to minimize (and measure) the ensemble spread. More generally, POPS-constrained loss minimization with our ensemble ansatz is potentially sub-optimal in that $\pi^{*}_{E}({\bm{\Theta}})$ chooses ${\bm{\Theta}}^{*}_{\mathbf{X}}$ to minimize the average prediction error across all other ${\mathbf{X}}^{\prime}\in{\mathcal{X}}$ but is not guaranteed to be a minimizer of (18). Exploring this tradeoff (multiple POPS occupancy vs global prediction error) would entail a combinatorial scheme which (21) avoids. One strategy, to be investigated in future work, would be to cluster the loss-minimizing pointwise fits, try to find POPS-intersections for each cluster, then replace all members of that cluster with a single POPS-intersecting fit. A further limitation of (21) is that many pointwise fits may be near-degenerate, raising the generalization error and, in general, underestimating the average deviation of any test point from the minimum loss prediction. However, the predicted max/min range of any test point is expected to be robust in the limit $N/P\to\infty$, as this is invariant to how we weight the ensemble.
Furthermore, even if it were possible to find the true minimizer $\pi^{*}_{\mathcal{G}}({\bm{\Theta}})$ of ${\mathcal{G}}_{0}$, it is not clear a priori that the resulting distribution of predicted test errors would be close to observed test errors. The covering constraint (15) only ensures that any minimizer of ${\mathcal{G}}_{0}$ should produce predictions which bound any test point, i.e. provide a max/min envelope on model errors. In practice, we will show that uniformly sampling from either the parameters or posterior predictions of the POPS-constrained loss minimizing ensemble yields very good results for a diverse set of linear regression problems, which we demonstrate in the remainder of this paper.
III.2 POPS-constrained ensemble for linear regression
@@
Linear models use feature functions $\mathcal{F}:{\mathcal{X}}\to{\mathcal{P}}\times{\mathcal{Y}}$ to map inputs ${\mathbf{X}}\in{\mathcal{X}}$ to features ${\mathbf{F}}({\mathbf{X}})\in{\mathcal{P}}\times{\mathcal{Y}}$, giving ${\mathcal{M}}({\mathbf{X}},{\bm{\Theta}})={\bm{\Theta}}\cdot{\mathbf{F}}({\mathbf{X}})$ with parameters ${\bm{\Theta}}\in{\mathcal{P}}$. Minimizing the PAC-Bayes bound (7) with a Gaussian prior $\pi_{0}({\bm{\Theta}})=\mathcal{N}({\bm{\Theta}}|{\bf 0},{\bm{\Sigma}}_{0})$ gives a minimizer ${\bm{\Theta}}^{*}_{\mathcal{L}}$ of
@@
We discuss how the epistemic covariance for small $N/P$ can be incorporated in a variational PAC-Bayes scheme in Appendix D. However, in the following, we concentrate on misspecification uncertainties in the limit $N/P\to\infty$, where the influence of $\bm{\Sigma}_{0}$ is negligible. For linear models the POPS-constraint is simply ${\bm{\Theta}}_{\mathbf{X}}\cdot{\mathbf{F}}({\mathbf{X}})\equiv{\mathcal{E}}({\mathbf{X}})$. The loss minimizing pointwise fits ${\bm{\Theta}}^{*}_{{\mathbf{X}}}$ can be evaluated analytically via the efficient rank-one updates
@@
where ${\rm h}({\mathbf{X}})=[{\mathbf{F}}({\mathbf{X}})]^{\top}{\bf A}{\mathbf{F}}({\mathbf{X}})$ is the leverage, an outlier measure closely related to the distance of (Mahalanobis, 1936). The minimum loss solution is the leverage-weighted centroid of pointwise optima, ${\bm{\Theta}}^{*}_{\mathcal{L}}=\langle{\rm h}({\mathbf{X}}){\bm{\Theta}}^{*}_{{\mathbf{X}}}\rangle_{\mathcal{X}}$. Equation (23) is our third main result, a scheme to incorporate misspecification uncertainty into least-squares linear regression for minimal overhead, requiring only $\mathcal{O}(N)$ inner products over feature vectors. Implementation in least squares linear regression is trivial, as shown in appendix F. We now investigate how our POPS-constrained ensemble performs in a variety of high-dimensional regression tasks.
IV Numerical Experiments
The supplementary material provides python/JaX code to reproduce all figures, with the exception of the large atomic machine learning datasets.
IV.1 Comparison with direct minimization of $\mathcal{G}$
For low dimensional problems, it is possible to approximately minimize an empirical generalization error $\mathcal{G}$ over discrete model ensembles, using a finite aleatoric uncertainty ${\bm{\Sigma}}_{\mathcal{Y}}$ for regularization. With increasing ${\bm{\Sigma}}_{\mathcal{Y}}$ it is clear that $\mathcal{G}\to\mathcal{L}$ and thus we expect the minimum loss solution $\pi_{\mathcal{L}}({\bm{\Theta}})$ to emerge. We also compare these appraoches to Bayesian ridge regression, corresponding to minimizing a first-order PAC-Bayes loss bound, and an ensemble approach from citepmasegosa2020learning, section C.3, which corresponds to minimizing a second-order PAC-Bayes loss bound, whose loss function, as for $\mathcal{G}_{E}$, also requires exponentiating model errors. At small ${\bm{\Sigma}}_{\mathcal{Y}}$ direct minimization of $\mathcal{G}_{E}$ and the Masegosa ensemble bound typically fails, requiring careful parameter tuning and remaining at low $P<6$ to avoid vanishing gradients or other numerical issues. The POPS-ensemble, in contrast, can easily sample $P=\mathcal{O}(1000)$ or higher dimensional data.
Figure 1 shows a simple example fitting a one dimensional quadratic polynomial to a sinusoidal function, with the constant, linear and quadratic features giving $P=3$ model parameters. However, the features are highly correlated, simplifying the regression task. We show results for ${\bm{\Sigma}}_{\mathcal{Y}}=1/3$, at the lower limit for which minimization converges for all methods. As expected from section II.3 Bayesian ridge regression significantly underestimates test errors even at the ±$\pm 3\sigma$ level, which as discussed above will worsen with increasing $N/P$. We find similar behaviour for the Masegosa ensemble, which we note is primarily designed for cases with finite aleatoric uncertainty. Direct minimization of $\mathcal{G}_{E}$ gives better bounding of test errors, though still with notable underestimates in high-density regions, though bounnds became tighter with lower ${\bm{\Sigma}}_{\mathcal{Y}}$ (whilst stable) However, the POPS-constrained ensemble gives almost perfect bounding of test errors, being stable for arbitrarly small ${\bm{\Sigma}}_{\mathcal{Y}}$ values.
@@ IV.1 Comparison with direct minimization of $\mathcal{G}$
Figure 1: Regression of a deterministic quadratic polynomial ($P=3$) model onto a sinusoidal ”simulation engine”, trained on $N=100$ points. Clockwise from top left: mean and $3\sigma$ confidence interval from Bayesian ridge regression , direct minimization of $\mathcal{G}_{E}$, regularized with $\Sigma_{\mathcal{Y}}=1$, the min/max range of the POPS-ensemble (23), and $\mathcal{G}_{E}$ with $\Sigma_{\mathcal{Y}}=0.1$.
We applied the same procedure to a model system with $P=6$ decorrelated features, unlike the polynomial example, to simulate more realistic datasets. The minimization routine could only treat a small range of $\Sigma_{\mathcal{Y}}$ values, which we normalize by the loss minimizer RMS error $\sigma_{MLE}$. Figure 2 shows the resulting parameter distributions. At large $\Sigma_{\mathcal{Y}}$ the minimizer quickly concentrates all ensemble members on the minimum loss solution, whilst at $\Sigma_{\mathcal{Y}}=\sigma_{MLE}$ the final distribution is very close to the POPS-ensemble. At lower values the final distribution was a little broader, but we attribute this largely to numerical instabilities as the final result is sensitive to initial conditions (unlike the first two cases). In addition, a minimizer initialized with the POPS-ensemble did not change appreciably. These albeit simple results give strong evidence that the POPS-ensemble is indeed a stable and highly efficient means to find approximate minimizers of ${\mathcal{G}}$ that would otherwise be intractable.
Figure 2: Regression of a $P=6$ linear model onto a cubic ”simulation engine”, trained on $N=500$ points. The panels show the culmulative distribution of each parameter across the ensemble (same size as $N$) for the loss minimizer (MLE), the POPS-ensemble (23) and from direct minimization of ${\mathcal{G}}$ with $\Sigma_{\mathcal{Y}}/\sigma_{MLE}=2,1,1/2$.
IV.2 Test errors in medium-dimensional linear regression

### v3

@@ II Analysis of the generalization error in the near-deterministic limit
for any $c>0$. Equation (15) is our second main result, a demonstration that any candidate minimizer of the generalization error must have probability mass in every POPS ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$, whilst the minimizing distribution concentrates on regions where multiple POPS intersect. As shown above, the minimum loss solution $\pi_{\mathcal{L}}({\bm{\Theta}})$ does not satisfy this requirement and thus ${\mathcal{G}}[\pi_{\mathcal{L}}]=\mathcal{O}(1/\epsilon^{2})$. Importantly, when (15) is satisfied, the predictive distribution $\rho_{\mathcal{M}}({\mathbf{Y}},{\mathbf{X}})$ covers every training data point, i.e. model predictions will envelope observations as $N/P\to\infty$. We now investigate two candidate distributions which satisfy (15), an $\mathcal{O}(N)$ ensemble ansatz and $\mathcal{O}(P)$ hypercube ansatz, finding approximate variational minima that allow for efficient deployment on high-dimensional datasets.
II.1 POPS-constrained ensemble and hypercube ansatz
The covering constraint (15) requires $\pi({\bm{\Theta}})$ to have mass in the POPS of each training point. This condition must be satisfied whilst concentrating mass in as small a region in ${\mathcal{P}}$ as possible to minimize ${\mathcal{G}}_{0}$. To achieve this, we first select a set of pointwise fits ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E}$ that are each POPS-constrained loss minimizers
${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E},\quad{\bm{\Theta}}^{*}_{\mathbf{X}}\equiv\arg\min_{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}L({\bm{\Theta}}_{\mathbf{X}}).$ (16)
These pointwise fits will be closely clustered around the minimum loss solution ${\bm{\Theta}}^{*}_{\mathcal{L}}$, as measured by the Fisher metric ${\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}$. In the supplementary materials Swinburne and Perez (2024) we explore a weighted ensemble ansatz
$\pi^{*}_{E}({\bm{\Theta}})=\langle{\rm w}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathbf{X}})\rangle,$ (17)
where ${\rm w}({\mathbf{X}})>0$. In particular, we show $\mathcal{T}^{*}_{E}$ naturally emerges when choosing pointwise fits ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$ to minimize the ansatz expected loss $\mathcal{L}[\pi_{E}]=\langle{\rm w}({\mathbf{X}})L({\bm{\Theta}}_{X})\rangle$.
The ensemble ansatz produces highly informative max/min bounds for any test point, providing an envelope for worst case errors, as we show in figure 1. However, even when employing an optimal reweighting ${\rm w}^{*}({\mathbf{X}})$ (see Swinburne and Perez (2024)), the ensemble ansatz systematically underestimates moments of the test error distribution, as shown in figure 2. In addition, both resampling and storage are $\mathcal{O}(N)$, which can become problematic for the $N/P\to\infty$ limit of interest.
For practical applications, the parameter distribution should ideally require at most $\mathcal{O}(P^{2})$ effort for storage and $\mathcal{O}(P)$ for resampling. We have found this can be achieved by finding the minimal hypercube $\mathcal{H}(\mathcal{T}^{*}_{E})\in\mathcal{P}$ which encompasses all members of the loss minimizing POPS-ensemble ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E}$. As $\mathcal{H}(\mathcal{T}^{*}_{E})$ intersects all POPS by construction, equation (15) is satisfied by the hypercube ansatz uniform over $\mathcal{H}(\mathcal{T}^{*}_{E})$, i.e.
$\pi^{*}_{\mathcal{H}}({\bm{\Theta}})=V^{-1}_{\mathcal{H}}\delta({\bm{\Theta}}\in\mathcal{H}(\mathcal{T}^{*}_{E})),$ (18)
where $V_{\mathcal{H}}=\int_{\mathcal{P}}\delta({\bm{\Theta}}\in\mathcal{H}(\mathcal{T}^{*}_{E})){\rm d}{\bm{\Theta}}$. Crucially, resampling $\pi_{\mathcal{H}}({\bm{\Theta}})$ requires only $\mathcal{O}(P)$ effort whilst giving a conservative estimate and bounding of test errors.
To determine $\mathcal{H}(\mathcal{T}^{*}_{E})$ we use singular value decomposition on the POPS-ensemble $\mathcal{T}^{*}_{E}$ to obtain the right eigenmatrix ${\bf V}\in\mathbb{R}^{R\times P}$, $R\leq P$, ${\bf V}{\bf V}^{\top}=\mathbb{I}_{R}$, requiring $\mathcal{O}(P^{3})$ effort for evaluation. Uniformly sampling the full max/min hypercube of the projected ensemble $\tilde{\mathcal{T}}^{*}_{E}=\{{\bf V}{\bm{\Theta}}|{\bm{\Theta}}\in{\mathcal{T}^{*}_{E}}\}$ returns samples $\tilde{{\bm{\Theta}}}^{\top}{\bf V}$ of $\pi^{*}_{\mathcal{H}}$ for $\mathcal{O}(R)$ effort and $\mathcal{O}(RP)$ storage. We find this efficient hypercube posterior to give excellent bounding and prediction of test errors for $P\in[3,1600]$, as shown in figures 2 and 3. An additional advantage is the simplicity of implementation for linear models; an example Python code is provided in the SM Swinburne and Perez (2024).
II.2 POPS-constrained loss minimization for linear models
@@
Both the ensemble and hypercube ansatz require POPS-constrained loss minimizing fits ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{P}_{\mathcal{M}}({\mathbf{X}})$. Practical application therefore requires an efficient means to perform this constrained minimization. Linear models use feature functions $\mathcal{F}:{\mathcal{X}}\to{\mathcal{P}}\times{\mathcal{Y}}$ to map inputs ${\mathbf{X}}\in{\mathcal{X}}$ to features ${\mathbf{F}}({\mathbf{X}})\in{\mathcal{P}}\times{\mathcal{Y}}$, giving ${\mathcal{M}}({\mathbf{X}},{\bm{\Theta}})={\bm{\Theta}}\cdot{\mathbf{F}}({\mathbf{X}})$ with parameters ${\bm{\Theta}}\in{\mathcal{P}}$. Minimizing the PAC-Bayes bound (21) with a Gaussian prior $\pi_{0}({\bm{\Theta}})=\mathcal{N}({\bm{\Theta}}|{\bf 0},{\bm{\Sigma}}_{0})$ gives a global loss minimizer ${\bm{\Theta}}^{*}_{\mathcal{L}}$ of
@@
The epistemic covariance for small $N/P$ can be incorporated in a variational PAC-Bayes scheme Swinburne and Perez (2024) that will be developed further in future work. However, in the following, we concentrate on misspecification uncertainties in the limit $N/P\to\infty$, where the influence of $\bm{\Sigma}_{0}$ is negligible. For linear models the POPS-constraint is simply ${\bm{\Theta}}_{\mathbf{X}}\cdot{\mathbf{F}}({\mathbf{X}})\equiv{\mathcal{E}}({\mathbf{X}})$. The loss minimizing pointwise fits ${\bm{\Theta}}^{*}_{{\mathbf{X}}}$ can be evaluated analytically via the efficient rank-one updates
@@
where ${\rm h}({\mathbf{X}})=[{\mathbf{F}}({\mathbf{X}})]^{\top}{\bf A}{\mathbf{F}}({\mathbf{X}})$ is the leverage, an outlier measure closely related to the distance of (Mahalanobis, 1936). The minimum loss solution is the leverage-weighted centroid of pointwise optima, ${\bm{\Theta}}^{*}_{\mathcal{L}}=\langle{\rm h}({\mathbf{X}}){\bm{\Theta}}^{*}_{{\mathbf{X}}}\rangle_{\mathcal{X}}$. Equation (20) is our third main result, a scheme to incorporate misspecification uncertainty into least-squares linear regression for minimal overhead, requiring only $\mathcal{O}(N)$ inner products over feature vectors, which can then be used to construct the ensemble or hypercube ansatz given above. A simple python implementation is provided in the SM Swinburne and Perez (2024).
III Numerical Experiments
III.1 Comparison to Bayesian ridge regression and minimization of $\mathcal{G}_{E}$
We applied Bayesian ridge regression and our POPS-ansatz to regress a quadratic polynomial against a sinusoidal function, the constant, linear and quadratic features giving $P=3$ model parameters. The SM Swinburne and Perez (2024) shows the same procedure for a model with $P=6$ independent features. For these simple models it is possible to approximately minimize the generalization error $\mathcal{G}_{E}$ of some discrete model ensemble, providing we regularize with a finite aleatoric uncertainty ${\bm{\Sigma}}_{\mathcal{Y}}=\sigma{\bm{\Sigma}}^{*}_{\mathcal{L}}$. With increasing $\sigma$ we have $\mathcal{G}\to\mathcal{L}$ and thus expect the ensemble to concentrate on ${\bm{\Theta}}^{*}_{\mathcal{L}}$, whilst in the small $\sigma$ regime of interest minimizing $\mathcal{G}_{E}$ becomes increasingly unstable. We found the same small $\sigma$ instability when applying an ensemble approach from Masegosa (2020) which targets a second-order PAC-Bayes loss bound, designed for misspecified probabilistic models where $\sigma$ can be large. These instabilities show existing misspecification-aware methods require careful hyperparameter tuning and low model dimension to avoid vanishing gradients or other numerical issues. Our POPS-ansatz, in contrast, can easily sample $P=\mathcal{O}(1000)$ or higher dimensional data.
Figure 1 shows results for $\sigma=1,1/3,1/6$, at the lower limit for which minimization converges. As expected from section I.4, Bayesian ridge regression significantly underestimates test errors even at the 99.7% confidence level (three standard deviations) which as discussed above will worsen with increasing $N/P$. The max/min range of the numerical ensemble shows similar behavior for $\sigma=1,1/3$ whilst at $\sigma=1/6$ test errors are bounded but significantly overestimated, which worsens as $\sigma$ decreases (whilst stable). This was also seen on higher dimensional datasets Swinburne and Perez (2024). Importantly, in all cases our POPS ansatz $\pi^{*}_{E}$ and $\pi^{*}_{\mathcal{H}}$ give excellent bounding of test errors and are stable in the near-deterministic limit $\sigma\to 0$. As expected, the hypercube ansatz $\pi^{*}_{\mathcal{H}}$ gave slightly more conservative bounds.
@@ III.1 Comparison to Bayesian ridge regression and minimization of $\mathcal{G}_{E}$
Figure 1: Regression of a deterministic quadratic polynomial ($P=3$) model onto a sinusoidal ”simulation engine”, trained on $N=100$ points. Top left: mean and $3\sigma$ interval from Bayesian ridge regression. All other plots show mean and max/min range. Top: POPS-ensemble and POPS-hypercube ansatz $\pi^{*}_{E}$ and $\pi^{*}_{\mathcal{H}}$. Bottom: numerically optimized $\mathcal{G}_{E}$ for a uniformly weighted $N$-ensemble, regularized with $\Sigma_{\mathcal{Y}}=\sigma\Sigma^{*}_{\mathcal{L}}$, for $\sigma=1,1/3,1/6$. Lower $\sigma$ values gave numerical instabilities.
III.2 Test errors in medium-dimensional linear regression